### v3

@@ 5 Adding accurate dynamical masses and radii to increase asteroseismic precision
Figure 8 depicts the complementarity among the methods of spectroscopy, asteroseismology, and binarity for stellar modelling. The top row illustrates the type of observations for each method: stellar spectra for atmospheric analysis, periodic brightness variations caused by stellar oscillations for asteroseismic analysis, and periodic dimming of the stellar light and radial velocity variations for binary analysis. The bottom row highlights the gain in precision and accuracy in the modelling as one combines spectroscopic and asteroseismic methods, at the same time benefiting from the accurate masses of the star from the binary dynamics. If the position of the star in the - Kiel diagram is only constrained by its atmospheric parameters inferred from spectroscopy, the typical error box (blue error box in the bottom left panel) is too large to conclude if the star will follow an evolutionary track of a 7 M intermediate-mass star with a white dwarf as the end-of-life product, or an 8 M star that ends its life as a neutron star. The effects of the initial metallicity (red versus black line) and of core boundary mixing (solid versus dashed line) on the evolutionary path of the star are too small to be distinguished with spectroscopic methods alone.
@@
Asteroseismology offers unique sensitivity to the physics in the deep stellar interior and is largely complementary to binary and atmospheric analyses in terms of the physical mechanisms these methods probe. Adding asteroseismic constraints delivers the smaller blue box instead of the grey one for a pure spectroscopic solution in the bottom middle panel. This asteroseismic addition improves the modelling substantially as it allows us to eliminate some of the highly improbably solutions, such as those based on the higher- and lower metal-content for the 7 M and 8 M models, respectively. Moreover, asteroseismic constraints on the amount and functional form of the near-core mixing is often tighter than suggested by the blue error box in the bottom middle panel in Fig. 8, such that the figure represents a pessimistic scenario. Ultimately, the additional model-independent constraint from binary dynamics decreases the error box to a fraction of the pure spectroscopic error box or the combined spectroscopic and asteroseismic error box (blue versus grey boxes in the bottom right panel). This way, one can characterise the physics of the binary pulsator with high accuracy.
@@
Finding binaries for which all three analysis methods depicted in Fig. 8 can be combined is a difficult task and such studies are therefore still scarce in the literature. The following criteria need to be met simultaneously to do the above-described analysis: (i) the object is an eclipsing or astrometric binary with high-quality photometric or astrometric data; (ii) the object is a spectroscopic double-lined binary such that the radial velocities of both binary components can be measured and their masses can be inferred; (iii) at least one of the components is a pulsator with time-series photometry of its brightness variations sufficiently long to achieve high-precision frequencies for sufficient modes; and (iv) at least some of the detected modes can be identified in terms of their . Even if these four conditions are met, it may still be impossible to determine which of the detected oscillation frequencies originate in which star, notably in binaries with components of almost equal mass.
@@
Various slowly rotating low-mass solar-like pulsators in eclipsing or astrometric binaries satisfy all four criteria. Moreover, because the mass and radius of a solar-like pulsator can be inferred directly from its oscillations through scaling relations when the effective temperature of the star is known (Kjeldsen & Bedding 1995), these asteroseismically inferred quantities can be confronted with the model-independent counterparts inferred from binary dynamics (e.g., Gaulme et al. 2013; Appourchaux et al. 2015). While asteroseismic measurements of mass and radius are in a good agreement with the dynamical values for unevolved solar-like pulsators (e.g., White et al. 2017), the asteroseismic masses and radii of red giants are systematically overestimated (e.g., Brogaard et al. 2016; Gaulme et al. 2016; Themeßl et al. 2018; Kallinger et al. 2018; Benbakoura et al. 2021).
@@
Intermediate-mass p-mode Sct pulsators in eclipsing binaries constitute another category of objects for applications, although mode identification remains a challenge for such moderate to fast rotators (cf. Sect. 4.3). For this reason, validation of stellar models with Sct pulsators in eclipsing binaries is to date primarily limited to the use of empirical relations between their oscillation properties and mean densities obtained from binary dynamics (e.g., García Hernández et al. 2015, 2017). Because many Sct pulsators are found in close Algol-type binary systems (a.k.a oEA stars; Mkrtichian et al. 2004, 2022) and p-mode oscillations reach their highest amplitudes in the outer layers of stars, oEA stars are often exploited to study binary interactions (see Sect. 6). As for high-mass p-mode pulsators, the discovery of 78 Cep pulsators in eclipsing binaries by Eze & Handler (submitted) in TESS data is extremely promising to calibrate binary supernova progenitors, of which some may be on their way to becoming gravitational wave emitters.
@@
The group of g-mode pulsators in eclipsing binaries is among the most promising in the context of validation and calibration of stellar models. The main reason lies in the effective probing power of the deep stellar interior facilitated by g modes, in particular the near-core regions (cf. Sect. 4.4). Binary and asteroseismic analyses become truly complementary in this particular regime and make strategies as in Fig. 8 optimally suited. To date, only a limited number of Dor, SPB stars, or white dwarf g-mode pulsators in eclipsing binaries with mode identification are known. Gaulme & Guzik (2019) presented a compilation of some 300 pulsating stars in eclipsing binaries observed with Kepler of which about a third have Dor pulsators. Southworth & Bowman (2022) and Southworth & Van Reeth (2022) reported the detection of 22 Dor, SPB, and Cep pulsators in eclipsing binaries observed with TESS. For now, none of those systems have identified modes, though.
@@
Sekaran et al. (2020) undertook a systematic search for unevolved g-mode pulsators in eclipsing binaries with period spacing patterns. Among the 93 detected eclipsing systems with a g-mode pulsating component, seven were found to exhibit period spacing patterns. In a follow-up study, Sekaran et al. (2021) investigated one of the detected systems in great detail. This concerns the slow rotator KIC 9850387, whose dynamical parameters are found to be in agreement with the parameters extracted from asteroseismic modelling provided that a high level of internal mixing for the pulsating component is used for its internal structure. The authors posited that this is a result of intrinsic non-tidal mixing mechanisms. The findings in Sekaran et al. (2021) reinforce the common conclusion in the literature from large samples of single-star pulsators and non-pulsating eclipsing binaries: stellar models require extra internal mixing compared to standard models to achieve a meaninful calibration. Similar studies of moderate to fast rotators with g modes in detached eclipsing binaries are not yet available, given that almost all such objects reveal tidal interactions as discussed in the next section.
@@
6 Tidal asteroseismology: the interplay between rotation, tides, and oscillations in binaries
@@
The space photometry revolution gave rise to a “zoo” of observed phenomena in Kepler and TESS data caused by tidal interactions in close binaries. Systems with free oscillation modes perturbed by tides, with tidally excited modes, with tidally tilted pulsation axes, with tidal deformation of the mode cavities, and with non-linear mode excitation due to tides have all been found (Guo 2021, for a review). In order to study the effect of linear tides in binaries with rotating components we need to solve Eq. (3), while additional terms as in Eq. (2) have to be included for non-linear tides.
@@
Tidally excited oscillations occur at exact multiples of the orbital frequency. They come in various flavours and may be caused by an equilibrium tide in case of a circular orbit and/or dynamical tides in eccentric orbits. An equilibrium tide is a static phenomenon occurring due to the balance between the pressure force and gravity and manifests itself as a geometrical deformation of one star due to the force exerted by the companion. Dynamical tides, on the other hand, cause a periodic acceleration due to in Eq. (3) as the components revolve around each other in their orbit (see the pedagogical paper by Fuller 2017). Tidal oscillations give rise to slow waves at low frequencies, which naturally occur in close binaries because the orbital, rotation, and g-mode periods are all roughly of similar order and reside in the left part of Fig. 3. On the other hand, self-excited oscillations may get perturbed due to both the equilibrium and dynamical tides, and such deformed oscillations are referred to as tidally perturbed oscillations. They may occur across the whole frequency axis in Fig. 3, including short-period p modes, when tidal forces are significant compared to the other forces at play in the system.
@@
Tidally excited oscillations represent the main mechanism of tidal dissipation in binaries composed of intermediate- or high-mass stars with a convective core and a radiative envelope (Zahn 2013; Alvan et al. 2013; Ogilvie 2014). In a nutshell, the mechanism acts as follows: (i) the tidal potential excites internal gravity waves in the deep stellar interior at the interface of the convective core and the radiative envelope of the star; (ii) these low-frequency travelling waves experience strong damping due to radiative diffusion (Zahn et al. 1997); (iii) the waves break in the low-density near-surface layers of the star, depositing their energy and angular momentum (Rogers et al. 2013), causing tidal heating and (pseudo-)synchronisation. Such tidal synchronisation is first initiated in the outer layers of the star and gradually proceeds inwards as more waves are being generated through tidal forcing and continue to dissipate. The dissipation is particularly efficient at so-called critical layers deep inside the star characterised by resonant waves having frequencies equal to the local co-rotation rate (Alvan et al. 2013).
@@
Tidally excited oscillations corresponding to standing waves occur at exact multiples of the orbital frequency. They do not suffer from strong damping and can therefore be observed in stars. Because the dominant component of the tide-generating potential is a spherical harmonic of , linear dynamical tides are expected to give rise to quadrupole modes. Some tidally excited oscillations can reach amplitudes exceeding a parametric instability threshold, in which case their mode energy will be transferred to daughter modes through non-linear mode coupling. The observational manifestation of such non-linear mode coupling is the occurrence of pairs or multiplets of daughter modes whose linear combination(s) satisfy resonance conditions (e.g., O’Leary & Burkart 2014; Manzoori 2020; Guo 2020, 2021; Guo et al. 2022). Moreover, some of the tidally excited oscillations can get into resonance locking with the orbital frequency of the binary. This happens when both frequencies evolve in concert. Such orbital resonant mode locking substantially increases the efficiency of tidal dissipation and orbital circularisation (e.g., Witte & Savonije 1999, 2001; Willems et al. 2003; Willems 2003; Burkart et al. 2012; Fuller & Lai 2012; Fuller et al. 2017; Fuller 2017; Zanazzi & Wu 2021; Kołaczek-Szymański & Różański 2023).
@@
It was recently also shown by Li et al. (2020a) and Fuller (2021) that “inverse tides” may occur in close binary systems hosting a g-mode pulsator. In this process, which occurs when the tidally forced mode amplitude approaches the mode’s saturation amplitude, a self-excited oscillation mode transfers (some of) its energy to the binary orbit. This may pump up the orbital eccentricity, create spin-orbit misalignment, and cause asynchronous rotation. This action of tidal dissipation is thus opposite to the standard view and rather pushes the state of the binary away from synchronised circular orbits. Another non-standard effect with respect to tidal theory in a two-body system was presented by Felce & Fuller (2023). They describe how the presence of a tertiary component may trap close binaries into a spin-orbit equilibrium known as a Cassini state. In this case, orbital precession due to the tertiary lies at the origin of the stable spin-orbit configuration. This mechanism has been invoked as an explanation for a minority of close binaries with stars of intermediate mass having extremely slow sub-synchronous rotation. Such a state can be understood by alterations of the spin-orbit misalignment angle and of the internal rotation rate caused by tidal dissipation due to inertial waves (Fuller 2021).
@@
Although the first (ground-based) observations of tidally excited oscillations in close binaries date from some two decades ago (Willems & Aerts 2002; Handler et al. 2002; Quaintrell et al. 2003), numerous detections had to await uninterrupted space-based photometric observations and gave rise to the great variety of observed behaviour we know today. Maceroni et al. (2009) reported the detection of tidally-excited oscillations in the eccentric binary HD 174884 from CoRoT data. A few years later, Welsh et al. (2011) presented spectacular tidally-excited oscillations at the 90 and 91 harmonic of the orbital frequency in HD 187091, also known as KOI-54. The light curve of KOI-54 is characterised by periodic brightness variations near periastron passage caused by geometrical deformation of the star as a response to the instantaneous tidal force. KOI-54 became the prototype of the class of eccentric binaries exhibiting periodic variability in its light curve similar to cardiograms of human heartbeats. These eccentric binaries are therefore also dubbed “heartbeat stars” (Thompson et al. 2012). Just like KOI-54 itself, the majority of heartbeat stars exhibit tidally excited oscillations (e.g., Hambleton et al. 2013, 2016; Kjurkchieva et al. 2016; Pablo et al. 2017; Hambleton et al. 2018; Guo et al. 2017, 2019; Cheng et al. 2020; Kołaczek-Szymański et al. 2021; Jayasinghe et al. 2021; Ou et al. 2021b, a; Wrona et al. 2022; Sharma et al. 2022; Wang et al. 2023). This makes them unique astrophysical laboratories to validate and improve tidal excitation theory (e.g., Burkart et al. 2012; Fuller & Lai 2012; Fuller 2017; O’Leary & Burkart 2014; Guo et al. 2022; Sun et al. 2023a). We refer the reader to Ogilvie (2014) and Guo (2021, and references therein) for comprehensive discussions on the theoretical and observational aspects of tidally excited oscillations in close binaries.
@@
Another distinct group of pulsators in (close) binary stars is the class of stars with tidally perturbed (or affected) oscillations. Tidal deformation in a close binary system affects the mode propagation cavities inside the components, as well as the alignment of the pulsation axes. While the effect of the tidal force on the eigenfrequencies of the star has been investigated to a reasonable level of detail with polytropic models (e.g., Saio 1981; Smeyers & Martens 1983; Martens & Smeyers 1986; Reyniers & Smeyers 2003b, a; Smeyers 2005), the consequences of the equilibrium tide on the mode eigenfunctions has been studied only recently (Fuller et al. 2020). While the former effect results in tidal splitting of a self-excited oscillation mode of a given spherical degree , the Fuller et al. (2020) formalism allows for tidal coupling between tidally split multiplets with different angular degree . The effect of such tidal mode coupling can be three-fold: (i) tidal alignment, where the axis of the oscillations gets aligned with the tidal axis rather than with the rotation axis of the star; (ii) tidal trapping, such that oscillation modes are confined to regions corresponding to the tidal pole or equator; (iii) tidal amplification, increasing flux perturbations in the tidal polar regions due to the propagation of acoustic waves close to the stellar surface.
@@
Observationally, tidally perturbed oscillations (which include both tidally split and trapped oscillations) manifest themselves as (a series of) modes that are offset from their adjacent orbital harmonics and are (quasi-)equidistantly spaced with the orbital frequency of the star (e.g., Balona 2018; Lee 2021a; Steindl et al. 2021; Van Reeth et al. 2022; Jennings et al. 2023). Some of the class members exhibit readily interpretable observed frequency patterns due to one of the above theoretical scenarios (e.g. some of the tidally tilted pulsators; Handler et al. 2020; Kurtz et al. 2020; Rappaport et al. 2021; Jayaraman et al. 2022; Kahraman Aliçavuş et al. 2022; Aliçavuş et al. 2023). Yet, others show either very complex or too sparse patterns to interpret (e.g., Bowman et al. 2019; Southworth et al. 2020; Jerzykiewicz et al. 2020; Southworth 2021; Southworth et al. 2021; Kálmán et al. 2022; Van Reeth et al. 2023; Johnston et al. 2023; Kálmán et al. 2023). Furthermore, in the specific case of eclipsing binaries, the geometrical effect of eclipse mapping due to particular obscuration of the visible disk of the pulsator by its companion (Reed et al. 2001, 2005) causes extra orbital phase-dependent variability of the mode amplitudes. Depending on the mode geometry, on the misalignment of the axis with respect to the rotation axis of the star, and on the geometry of the eclipse, the amplitude modulation can manifest itself as a complex frequency pattern in the Fourier transform of the light curve as in the case of U Gru (Johnston et al. 2023).
@@
Recent advances in asteroseismology of (close) binary stars can hardly be overestimated, both from the theoretical and observational point of view. Major efforts are undertaken to develop general enough tidal oscillation theory as a pertinent need to explain the zoo of observed phenomena resulting from the interplay between tides and oscillations brought by the high-precision space photometry. The novel theories mentioned above are powerful in terms of the effects due to the tidal, Lorentz, Coriolis, and centrifugal forces acting on a rotating pulsating star in a close binary or triple system, taking into account misalignments between the axes of apsides and rotation (Fuller et al. 2020; Fuller 2021; Sun et al. 2023a).

### v4

@@ 5 Adding accurate dynamical masses and radii to increase asteroseismic precision
Figure 8 depicts the complementarity among the methods of spectroscopy, asteroseismology, and binarity for stellar modelling. The top row illustrates the type of observations for each method: stellar spectra for atmospheric analysis, periodic brightness variations caused by stellar oscillations for asteroseismic analysis, and periodic dimming of the stellar light and radial velocity variations for binary analysis. The bottom row highlights the gain in precision and accuracy in the modelling as the spectroscopic and asteroseismic methods are combined, at the same time benefiting from the accurate masses of the star from the binary dynamics. If the position of the star in the - Kiel diagram is only constrained by its atmospheric parameters inferred from spectroscopy, the typical error box (blue error box in the bottom left panel) is too large to conclude whether the star will follow an evolutionary track of a 7 M⊙ intermediate-mass star with a white dwarf as the end-of-life product, or the one of an 8 M⊙ star that ends its life as a neutron star. The effects of the initial metallicity (red versus black line) and of core boundary mixing (solid versus dashed line) on the evolutionary path of the star are too small to be distinguished with spectroscopic methods alone.
@@
Asteroseismology offers unique sensitivity to the physics in the deep stellar interior and is largely complementary to binary and atmospheric analyses in terms of the physical mechanisms these methods probe. Adding asteroseismic constraints delivers the smaller blue box instead of the grey one for a pure spectroscopic solution in the bottom middle panel. This asteroseismic addition improves the modelling substantially as it allows us to eliminate some of the highly improbably solutions, such as those based on the higher- and lower metal content for the 7 M⊙ and 8 M⊙ models, respectively. Moreover, asteroseismic constraints on the amount and functional form of the near-core mixing is often tighter than suggested by the blue error box in the bottom middle panel in Fig. 8, such that the figure represents a pessimistic scenario. Ultimately, the additional model-independent constraint from binary dynamics decreases the error box to a fraction of the pure spectroscopic error box or the combined spectroscopic and asteroseismic error box (blue versus grey boxes in the bottom right panel). This way, one can characterise the physics of the binary pulsator with high accuracy.
@@
Finding binaries for which all three analysis methods depicted in Fig. 8 can be combined is a difficult task, and such studies are therefore still scarce in the literature. The following criteria need to be met simultaneously to do the above-described analysis: (i) the object is an eclipsing or astrometric binary with high-quality photometric or astrometric data; (ii) the object is a spectroscopic double-lined binary such that the radial velocities of both binary components can be measured and their masses can be inferred; (iii) at least one of the components is a pulsator with time series photometry of its brightness variations sufficiently long to achieve high-precision frequencies for sufficient modes; and (iv) at least some of the detected modes can be identified in terms of their . Even if these four conditions are met, it may still be impossible to determine which of the detected oscillation frequencies originates in which star, notably in binaries with components of almost equal mass.
@@
Various slow-rotating low-mass solar-like pulsators in eclipsing or astrometric binaries satisfy all four criteria. Moreover, because the mass and radius of a solar-like pulsator can be inferred directly from its oscillations through scaling relations when the effective temperature of the star is known (Kjeldsen & Bedding 1995), these asteroseismically inferred quantities can be confronted with the model-independent counterparts inferred from binary dynamics (e.g. Gaulme et al. 2013; Appourchaux et al. 2015). While asteroseismic measurements of mass and radius are in a good agreement with the dynamical values for unevolved solar-like pulsators (e.g. White et al. 2017), the asteroseismic masses and radii of red giants are systematically overestimated (e.g. Brogaard et al. 2016; Gaulme et al. 2016; Themeßl et al. 2018; Kallinger et al. 2018; Benbakoura et al. 2021).
@@
Intermediate-mass p-mode Sct pulsators in eclipsing binaries constitute another category of objects for applications, although mode identification remains a challenge for such moderate to fast rotators (see Sect. 4.3). For this reason, validation of stellar models with Sct pulsators in eclipsing binaries is currently primarily limited to the use of empirical relations between their oscillation properties and mean densities obtained from binary dynamics (e.g. García Hernández et al. 2015, 2017). Because many Sct pulsators are found in close Algol-type binary systems (also known as oEA stars; Mkrtichian et al. 2004, 2022) and p-mode oscillations reach their highest amplitudes in the outer layers of stars, oEA stars are often exploited to study binary interactions (see Sect. 6). As for high-mass p-mode pulsators, the discovery of 78 Cep pulsators in eclipsing binaries by Eze & Handler (2024) in the TESS data is extremely promising to calibrate binary supernova progenitors, of which some may be on their way to becoming gravitational wave emitters.
@@
The group of g-mode pulsators in eclipsing binaries is among the most promising in the context of validation and calibration of stellar models. The main reason lies in the effective probing power of the deep stellar interior facilitated by g modes, in particular the near-core regions (see Sect. 4.4). Binary and asteroseismic analyses become truly complementary in this particular regime and make strategies shown in Fig. 8 optimally suitable. Currently, only a limited number of Dor, SPB stars, or white dwarf g-mode pulsators in eclipsing binaries with mode identification are known. Gaulme & Guzik (2019) presented a compilation of 300 pulsating stars in eclipsing binaries observed with Kepler of which about one-third have Dor pulsators. Among such pulsators found by Li et al. (2020a) is the remarkable quadruply eclipsing heptuple system KIC 4150611 (Kemp et al. 2024), whose primary exhibits a long series of prograde dipole gravito-inertial modes yet to be modelled asteroseismically. Further, Southworth & Bowman (2022) and Southworth & Van Reeth (2022) reported the detection of 22 Dor, SPB, and Cep pulsators in eclipsing binaries observed with TESS. For now, however, none of these systems have identified modes.
@@
Sekaran et al. (2020) undertook a systematic search for unevolved g-mode pulsators in eclipsing binaries with period spacing patterns. Among the 93 detected eclipsing systems with a g-mode pulsating component, seven were found to exhibit period spacing patterns. In a follow-up study, Sekaran et al. (2021) investigated one of the detected systems in great detail. This deals with the slow rotator KIC 9850387, whose dynamical parameters are found to be in agreement with the parameters extracted from asteroseismic modelling provided that a high level of internal mixing for the pulsating component is used for its internal structure. The authors proposed that this is a result of intrinsic non-tidal mixing mechanisms. The findings in Sekaran et al. (2021) reinforce the common conclusion in the literature from large samples of single-star pulsators and non-pulsating eclipsing binaries: stellar models require extra internal mixing compared to standard models to achieve a meaninful calibration. Similar studies of moderate to fast rotators with g modes in detached eclipsing binaries are not yet available, given that almost all such objects reveal tidal interactions, as discussed in the next section.
@@
6 Tidal asteroseismology: Interplay between rotation, tides, and oscillations in binaries
@@
The space photometry revolution gave rise to a ‘zoo’ of observed phenomena in Kepler and TESS data caused by tidal interactions in close binaries. Systems with free oscillation modes perturbed by tides, with tidally excited modes, with tidally tilted pulsation axes, with tidal deformation of the mode cavities, and with non-linear mode excitation due to tides have all been found (Guo 2021, for a review). In order to study the effect of linear tides in binaries with rotating components we need to solve Eq. (3), while additional terms (as in Eq. (2)) have to be included for non-linear tides.
@@
Tidally excited oscillations occur at exact multiples of the orbital frequency. They come in various flavours and may be caused by an equilibrium tide in case of a circular orbit and/or dynamical tides in eccentric orbits. An equilibrium tide is a static phenomenon occurring due to the balance between the pressure force and gravity and manifests itself as a geometrical deformation of one star due to the force exerted by the companion. Dynamical tides, on the other hand, cause a periodic acceleration due to in Eq. (3) as the components revolve around each other in their orbit (see the pedagogical paper by Fuller 2017). Tidal oscillations give rise to slow waves at low frequencies, which naturally occur in close binaries because the orbital, rotation, and g-mode periods are all of similar order and reside in the left part of Fig. 3. On the other hand, self-excited oscillations may get perturbed due to both the equilibrium and dynamical tides, and such deformed oscillations are referred to as tidally perturbed oscillations. They may occur across the whole frequency axis in Fig. 3, including short-period p modes, when tidal forces are significant compared to the other forces at play in the system.
@@
Tidally excited oscillations represent the main mechanism of tidal dissipation in binaries composed of intermediate- or high-mass stars with a convective core and a radiative envelope (Zahn 2013; Alvan et al. 2013; Ogilvie 2014). In a nutshell, the mechanism acts as follows: (i) the tidal potential excites internal gravity waves in the deep stellar interior at the interface of the convective core and the radiative envelope of the star; (ii) these low-frequency travelling waves experience strong damping due to radiative diffusion (Zahn et al. 1997); and (iii) the waves break in the low-density near-surface layers of the star, depositing their energy and angular momentum (Rogers et al. 2013), causing tidal heating and (pseudo-)synchronisation. Such tidal synchronisation is first initiated in the outer layers of the star and gradually proceeds inwards as more waves are being generated through tidal forcing and continue to dissipate. The dissipation is particularly efficient at the critical layers deep inside the star characterised by resonant waves having frequencies equal to the local co-rotation rate (Alvan et al. 2013).
@@
Tidally excited oscillations corresponding to standing waves occur at exact multiples of the orbital frequency. They do not suffer from strong damping, and can therefore be observed in stars. Because the dominant component of the tide-generating potential is a spherical harmonic of , linear dynamical tides are expected to give rise to quadrupole modes. Some tidally excited oscillations can reach amplitudes exceeding a parametric instability threshold, in which case their mode energy will be transferred to daughter modes through non-linear mode coupling. The observational manifestation of such non-linear mode coupling is the occurrence of pairs or multiplets of daughter modes whose linear combination(s) satisfy resonance conditions (e.g. O’Leary & Burkart 2014; Manzoori 2020; Guo 2020, 2021; Guo et al. 2022). Moreover, some of the tidally excited oscillations can enter into resonance, locking with the orbital frequency of the binary. This happens when the two frequencies evolve in concert. Such orbital resonant mode locking substantially increases the efficiency of tidal dissipation and orbital circularisation (e.g. Witte & Savonije 1999, 2001; Willems et al. 2003; Willems 2003; Burkart et al. 2012; Fuller & Lai 2012; Fuller et al. 2017; Fuller 2017; Zanazzi & Wu 2021; Kołaczek-Szymański & Różański 2023).
@@
It was recently also shown by Li et al. (2020a) and Fuller (2021) that ‘inverse tides’ may occur in close binary systems hosting a g-mode pulsator. In this process, which occurs when the tidally forced mode amplitude approaches the mode’s saturation amplitude, a self-excited oscillation mode transfers (some of) its energy to the binary orbit. This may pump up the orbital eccentricity, create spin-orbit misalignment, and cause asynchronous rotation. This action of tidal dissipation is thus opposite to the standard view and rather pushes the state of the binary away from synchronised circular orbits. Another non-standard effect with respect to tidal theory in a two-body system was presented by Felce & Fuller (2023). They describe how the presence of a tertiary component may trap close binaries into a spin-orbit equilibrium known as a Cassini state. In this case, orbital precession due to the tertiary lies at the origin of the stable spin-orbit configuration. This mechanism has been invoked as an explanation for a minority of close binaries with stars of intermediate mass having extremely slow sub-synchronous rotation. Such a state can be understood by alterations of the spin-orbit misalignment angle and of the internal rotation rate caused by tidal dissipation due to inertial waves (Fuller 2021).
@@
Although the first (ground-based) observations of tidally excited oscillations in close binaries date from two decades ago (Willems & Aerts 2002; Handler et al. 2002; Quaintrell et al. 2003), numerous detections had to await uninterrupted space-based photometric observations and gave rise to the great variety of observed behaviour we know today. Maceroni et al. (2009) reported the detection of tidally excited oscillations in the eccentric binary HD 174884 from CoRoT data. A few years later, Welsh et al. (2011) presented spectacular tidally excited oscillations at the 90 and 91 harmonic of the orbital frequency in HD 187091, also known as KOI-54. The light curve of KOI-54 is characterised by periodic brightness variations near periastron passage caused by geometrical deformation of the star as a response to the instantaneous tidal force. KOI-54 became the prototype of the class of eccentric binaries exhibiting periodic variability in its light curve similar to cardiograms of human heartbeats. These eccentric binaries are therefore also dubbed ‘heartbeat stars’ (Thompson et al. 2012). Just like KOI-54 itself, the majority of heartbeat stars exhibit tidally excited oscillations (e.g. Hambleton et al. 2013, 2016; Kjurkchieva et al. 2016; Pablo et al. 2017; Hambleton et al. 2018; Guo et al. 2017, 2019; Cheng et al. 2020; Kołaczek-Szymański et al. 2021; Jayasinghe et al. 2021; Ou et al. 2021b, a; Wrona et al. 2022; Sharma et al. 2022; Wang et al. 2023). This makes them unique astrophysical laboratories to validate and improve tidal excitation theory (e.g. Burkart et al. 2012; Fuller & Lai 2012; Fuller 2017; O’Leary & Burkart 2014; Guo et al. 2022; Sun et al. 2023a). We refer to Ogilvie (2014) and Guo (2021, and references therein) for comprehensive discussions on the theoretical and observational aspects of tidally excited oscillations in close binaries.
@@
Another distinct group of pulsators in close binary stars is the class of stars with tidally perturbed oscillations, which means that the star’s free oscillations are affected by the presence of a companion. Tidal deformation in a close binary system affects the mode propagation cavities inside the components, as well as the alignment of the pulsation axes. While the effect of the tidal force on the eigenfrequencies of the star has been investigated to a reasonable level of detail with polytropic models (e.g. Saio 1981; Smeyers & Martens 1983; Martens & Smeyers 1986; Reyniers & Smeyers 2003b, a; Smeyers 2005), the consequences of the equilibrium tide on the mode eigenfunctions has been studied only recently (Fuller et al. 2020). While the former effect results in tidal splitting of a self-excited oscillation mode of a given spherical degree , the Fuller et al. (2020) formalism allows tidal coupling between tidally split multiplets with different angular degree . The effect of such tidal mode coupling can be three-fold: (i) tidal alignment, where the axis of the oscillations gets aligned with the tidal axis rather than with the rotation axis of the star; (ii) tidal trapping, such that oscillation modes are confined to regions corresponding to the tidal pole or equator; and (iii) tidal amplification, increasing flux perturbations in the tidal polar regions due to the propagation of acoustic waves close to the stellar surface.
@@
Observationally, tidally perturbed oscillations (which include both tidally split and trapped oscillations) manifest themselves as (a series of) modes that are offset from their adjacent orbital harmonics and are (quasi-)equidistantly spaced with the orbital frequency of the star (e.g. Balona 2018; Lee 2021a; Steindl et al. 2021; Van Reeth et al. 2022; Jennings et al. 2024). Some of the class members exhibit readily interpretable observed frequency patterns due to one of the above theoretical scenarios (e.g. some of the tidally tilted pulsators; Handler et al. 2020; Kurtz et al. 2020; Rappaport et al. 2021; Jayaraman et al. 2022; Kahraman Aliçavuş et al. 2022; Aliçavuş et al. 2023). However, others show patterns that are either very complex or too sparse to interpret (e.g. Bowman et al. 2019; Southworth et al. 2020; Jerzykiewicz et al. 2020; Southworth 2021; Southworth et al. 2021; Kálmán et al. 2022; Van Reeth et al. 2023; Johnston et al. 2023; Kálmán et al. 2023). Furthermore, in the specific case of eclipsing binaries, the geometrical effect of eclipse mapping due to particular obscuration of the visible disk of the pulsator by its companion (Reed et al. 2001, 2005) causes extra orbital phase-dependent variability of the mode amplitudes. Depending on the mode geometry, on the misalignment of the axis with respect to the rotation axis of the star, and on the geometry of the eclipse, the amplitude modulation can manifest itself as a complex frequency pattern in the Fourier transform of the light curve, as in the case of U Gru (Johnston et al. 2023).
@@
Recent advances in asteroseismology of (close) binary stars can hardly be overestimated, both from the theoretical and observational point of view. Major efforts are undertaken to develop a general enough tidal oscillation theory as a pertinent need to explain the zoo of observed phenomena resulting from the interplay between tides and oscillations brought by the high-precision space photometry. The novel theories mentioned above are powerful in terms of the effects due to the tidal, Lorentz, Coriolis, and centrifugal forces acting on a rotating pulsating star in a close binary or triple system, taking into account misalignments between the axes of apsides and rotation (Fuller et al. 2020; Fuller 2021; Sun et al. 2023a).